The first term in a sequence of number is 2. Each even-numbered term is 3 more than the previous term and each odd-numbered term, excluding the first, is –1 times the previous term. What is the 45 th term of the sequence?
Answer by sudhanshu_kmr(1152)


Sequence is 2, (2+3) , (2+3)-1, (2+3)-1+3 , ................
i.e 2, 5, 4, 7, 6, 9, 8, 11............
here even position numbers are 2, 4, 6, 8, 10.........

it is in Arithmetic progression and difference = 2

odd position numbers are 5, 7, 9, 11............

also arithmetic progression , difference = 2
45th term would be 23th term of odd position numbers.
for nth term of a Arithmetic progression = a + (n-1) d

= 5 + (23 -1)2

= 5 + 22 * 2

= 49
so, required number is 49